# Tomoya Shiota,∗,†,a,b Kenji Ishihara,†,b and Wataru Mizukami∗,a,b

### v2

@@ 3.2 DFT verification of the LSE-based predictions
The accuracy of the adsorption energy predictions was verified using an NNP and the LSE with DFT calculations while seeking to increase the prediction accuracy. In our computing environment, the computation time for evaluating the adsorption energy via DFT was approximately $10^{3}$ times that of the NNP, as indicated by Table 1. Therefore, we randomly selected one of the 20 structures of $\text{HEA}_{201}$ discussed above and evaluated the adsorption energies for 19 sites on the (111) plane, as shown on the right side of Figure 4. The selected $\text{HEA}_{201}$ was $\text{Ir}_{40}\text{Pd}_{40}\text{Pt}_{41}\text{Rh}_{40}\text{Ru}_{40}$. First, the prediction accuracy of the adsorption energies between the prediction based on the LSE and direct NNP calculation was compared with that of the DFT calculations. The RMSE of the adsorption energy for all sites obtained via the direct NNP calculation was 0.445 eV. The RMSE value of the prediction based on the LSE was 0.425 eV, corresponding to a slightly higher accuracy (0.020 eV) compared with the direct evaluation via the NNP. Thus, the accuracy of the prediction model using the LSE was close to that of the NNP. However, comparison of the adsorption energy values from the DFT-parametrized LSE (I-LSE) model with those from the direct DFT calculations revealed a significant improvement in the prediction performance, with an RMSE of 0.234 eV, representing a 50% reduction in error. This outcome demonstrates the potential of the LSE to improve the adsorption energy prediction performance.
@@ 4 Discussion
Finally, we discuss potential applications of the proposed method for designing chemical reactions for catalyst and device development. Previous studies examined the role of atomic energy in increasing the accuracy and efficiency of NNPs and validated atomic-energy mapping results 60, 64, 65, 66, 67. Recently, studies have attempted to gain chemical insight from atomic energies 68. However, to the best of our knowledge, the proposed LSE prediction method is the first to demonstrate that atomic energy can serve as a descriptor for the efficient evaluation of catalytic properties. Moreover, in contrast to atomic energies, which are absolute quantities, LSEs are relative quantities and are thus expected to be less sensitive to differences in computational methods, such as the treatment of basis functions and inner-shell electrons. The proposed approach enhances chemical reaction design—a crucial component of machine learning-based material design—which has garnered significant attention in the scientific community over the past few years due to its promise to accelerate the discovery and development of useful materials.69, 70.
@@ 5 Conclusion
Building upon this work, future research can extend the application of the LSE descriptor to other molecular species and reaction systems, validating and enhancing its predictive accuracy across a broader spectrum of catalytic processes. Integrating the LSE descriptor with advanced machine learning algorithms could facilitate large-scale screening of HEA compositions, accelerating the discovery of optimal catalysts for specific reactions. Additionally, applying the LSE framework to other material systems—such as high-entropy nitrides, oxides, and carbides—could further expand its impact on materials design.
In this study, we assumed that the HEA NPs have random elemental configurations following a uniform distribution. However, short-range order is also important in high-entropy alloys71, 72, 73. By combining our method with techniques for quantifying short-range order using E(3)-equivariant graph neural networks 74, we can potentially obtain more accurate distributions of reactivity. This integration would allow us to capture the effects of atomic arrangements more precisely, leading to improved predictions of catalytic properties.
Moreover, this approach holds the potential to consider more realistic environmental conditions, including the behavior of catalysts in solution or under operational settings. Incorporating factors such as solvent effects, temperature, and pressure into the screening process would enhance the relevance and applicability of the predictions, leading to more effective and practical catalyst designs.
@@ 6.1 Modeling
We modeled CO adsorption on the NPs, i.e., CO/$\mathrm{M}_{\mathrm{n}}$, using an atomic simulation environment (ASE) 75, 76. The initial lattice constants (LCs) of $\mathrm{M}_{\mathrm{n}}$ were determined via bulk calculations. For $\mathrm{HEA}_{\mathrm{201}}$, the largest bulk was used as the initial LC. A 15-Å vacuum region was inserted in all supercells to minimize cell-to-cell interactions. The initial structures of $\mathrm{CO/M}_{\mathrm{n}}$ and $\mathrm{CO/HEA}_{\mathrm{201}}$ were derived by placing CO on the on-top sites of the optimized structures of $\mathrm{M}_{\mathrm{n}}$ and $\mathrm{HEA}_{\mathrm{201}}$, respectively, with the distance between the C atom and the adsorption-site metal atom M set as 2 Å.
@@ 6.2 NNP calculations
The NNP used was M3GNet, which is a pretrained model based on a graph neural network capable of computing periodic tables 36. The crystal structures of the bulk $\mathrm{M=Ir,Pd,Pt,Rh,Ru}$ were assumed to be fcc. Although $\mathrm{Ru}$ exhibits a hexagonal close-packed form at room temperature, $\mathrm{Ru}$ with an fcc structure can be created using NPs 77. The atomic energy of the bulk fcc metal M was determined using the energy corresponding to the minimum value obtained when varying the LC of the material at intervals of approximately 0.01 $\mathrm{\AA}$. The resulting atomic-energy values and corresponding LCs are presented in Table 2. The Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm was used for structural optimizations, with a maximum step of 0.005 Å. In the CO/M and CO/HEA structural optimization calculations, the M and HEA structures were fixed as stable isolated systems. Only the CO and adsorption-site atoms were relaxed, with the constraint that the adsorbed molecules occupy the top site. Structural optimizations were performed until the force acting on each atom was 0.001 eV/Å.
@@ 6.3 DFT calculations
The DFT calculations were performed using the Vienna Ab initio Simulation Package (VASP), version 5.4.4 78, 79, 80. We employed the PBE generalized gradient approximation functional as the exchange-correlation functional81. The core electrons were treated using the projector-augmented wave method 82, 83. Electronic structures were optimized using the blocked Davidson iteration scheme within a spin-restricted approximation. The cutoff energy for the plane-wave functions was set as 400 eV. The ion geometries were optimized using a conjugate-gradient algorithm with a convergence threshold of 0.01 eV/Å.
@@ 6.4 Nonlinear regressions
KRRs were executed using scikit-learn version 1.2.2 84. A Gaussian kernel was selected as the kernel function of the KRR. The hyperparameters for each model were optimized over 100 iterations using a randomized search.
QCL regressions were implemented using scikit-qulacs version 0.5.085. Qulacs version 0.5.6 was used as the quantum-circuit simulator86. The number of qubits in the QCL regression model was 4. The number of iterations of the parameterized variational quantum circuit of the model corresponding to the weights of the neural network was 6. The timestep for the time evolution operator in parameterized variational quantum circuits was set as 0.5. The BFGS algorithm was used to update the parameters of the QCL model.
@@
Table 3: Parameters $\alpha$ and $\beta$ of the regression lines between the CO adsorption energy and the local surface energy (LSE) for monometallic systems, obtained using NNP and DFT methods. The values correspond to Figures 2 and 7, respectively, where NNP and DFT in parentheses indicate the method used to obtain these parameters.
@@
Fig.  6: Correlation between the adsorption energies of a CO molecule on monometallic NPs calculated directly using the NNP and those predicted through regression. Circles, squares, and triangles at each datapoint represent the facet, edge, and corner CO adsorption sites, respectively.
Fig.  7: Correlation between adsorption energies of a CO molecule on monometallic NPs calculated directly via DFT and the LSEs corresponding to each adsorption site. The CO adsorption energies for the 38- and 79-atom NPs, as well as the datapoints corresponding to the hollow sites that converged on the Pd NPs, are shown as open markers, as they were not included in the linear regression data points. Circles, squares, and triangles at each datapoint represent the facet, edge, and corner CO adsorption sites, respectively.
Fig.  8: (a) Comparison and (b) correlation between the adsorption energy predicted via LSE and the adsorption energy of $\mathrm{CO/HEA}_{201}$ directly calculated using the NNP. We limit the comparison to 2440 sites, encompassing all on-top sites across the 20 structures. For each element (Ir, Pd, Pt, Rh, Ru) for 14 patterns of IrPdPtRhRu $\text{HEA}_{201}$ NPs, the energies are corrected using nonlinear regression models based on (c) KRR and (d) QCL regression. These values are plotted against values directly obtained using the NNP, which serve as the standard for accuracy.
@@
The data and code required to reproduce the figures and tables are publicly accessible on GitHub https://github.com/TShiotaSS/lse. NP datasets are available at DOI: https://figshare.com/articles/dataset/Nano_particle_structures/26973409. We modified the code to extract the atomic energies from the pretrained M3GNet model on Github at https://github.com/materialsvirtuallab/m3gnet, which can be found at https://github.com/TShiotaSS/lse/tree/main/scripts/m3gnet_each_atom_energy. The implementation for QCL regression used in this study is available at https://github.com/Qulacs-Osaka/scikit-qulacs/tree/main.
@@
Here, we summarize the parameters of the regression model for adsorption-energy prediction using the LSE descriptor as well as the results of the model accuracy validation. Figure 6 shows the parity plot of the training data and prediction results for the LSE linear regression model presented in Figure 2. The RMSE for predicting the training data was 0.035 eV. Figure 7 presents the regression results for the linear regression model between the CO adsorption energies on monometallic NPs obtained via DFT calculations and the LSE. Two datapoints corresponding to the hollow sites that converged on the facets of Pd in 38- and 79-atom monometallic NPs were excluded from the regression model. For small NPs, the entire system tended to exhibit molecular characteristics upon adsorption, making it difficult for a simple regression model to capture these effects . However, for larger NPs, a linear relationship with the LSE, similar to that observed for the NNP, was confirmed. The parameters of the regression models constructed using the adsorption energies from both the NNP and DFT are presented in Table 3. Figure 8 presents a parity plot of the prediction results for the data not included in the training data for the nonlinear regression model of HEA NP adsorption energies evaluated directly using the NNP. The prediction accuracies were 0.0580 and 0.0579 eV, respectively, indicating that, as demonstrated in this study, the regression models using KRR and QCL had comparable accuracy.

### v3

@@ 3.2 DFT verification of the LSE-based predictions
The accuracy of the adsorption energy predictions was verified using an NNP and the LSE with DFT calculations while seeking to increase the prediction accuracy. In our computing environment, the computation time for evaluating the adsorption energy via DFT was approximately $10^{3}$ times that of the NNP, as indicated by Table 1. Therefore, we randomly selected one of the 20 structures of $\text{HEA}_{201}$ discussed above and evaluated the adsorption energies for 19 sites on the (111) plane, as shown on the right side of Figure 5. The selected $\text{HEA}_{201}$ was $\text{Ir}_{40}\text{Pd}_{40}\text{Pt}_{41}\text{Rh}_{40}\text{Ru}_{40}$. First, the prediction accuracy of the adsorption energies between the prediction based on the LSE and direct NNP calculation was compared with that of the DFT calculations. The RMSE of the adsorption energy for all sites obtained via the direct NNP calculation was 0.445 eV. The RMSE value of the prediction based on the LSE was 0.425 eV, corresponding to a slightly higher accuracy (0.020 eV) compared with the direct evaluation via the NNP. Thus, the accuracy of the prediction model using the LSE was close to that of the NNP. Although the RMSE of 0.445 eV for direct NNP calculations may appear large, the M3GNet NNP systematically overestimates CO adsorption energies, with a mean error (ME) of 0.388 eV relative to DFT as shown in Figure 5. A similar overestimation trend is observed for monometallic NPs, where the RMSE and ME against DFT are 0.439 eV and 0.379 eV, respectively. As a result, the LSE-based prediction also displays a systematic overestimation of the CO adsorption energy in comparison with DFT. Notably, the adsorption energy range for the 19 HEA sites spans from 1.329 eV to 1.567 eV (by DFT and NNP, respectively), indicating a broad distribution of possible adsorption energies. Given this wide range, a constant shift between NNP and DFT remains comparatively tolerable for high-throughput screening of HEA catalysts.
While the M3GNet NNP has difficulty quantitatively describing CO-adsorbed states, it effectively captures the atomic-level stability of the NPs. This is supported by the small RMSD (0.09 Å) between the DFT-relaxed and NNP-relaxed HEA NP structures, as well as the LSE RMSE of 0.026 eV for the 19 target sites, indicating that M3GNet—trained on diverse crystal environments—can reliably describe complex HEA NP configurations despite their absence from its original training set. Consequently, by combining DFT-derived adsorption energies for monometallic NPs with M3GNet-derived LSE (i.e., the I-LSE method), the prediction error is halved to an RMSE of 0.234 eV. This improvement arises because the LSEs are calculated from the NPs before adsorption, thereby avoiding the systematic overestimation of adsorption energy found in direct NNP calculations, while retaining the efficiency of an LSE-based framework.
@@ 3.3 Distribution of the Adsorption Energy on HEA NPs
3.4 Universality of the LSE-based method
Fig.  7: Universality of the LSE-based prediction model for CO adsorption energy across different HEA nanoparticle (NP) shapes. (a) Predicted adsorption energies from the NNP for all ontop sites on a regular octahedron-shaped HEA146, plotted as a function of LSE. (b) Corresponding predictions for an icosahedron-shaped HEA147. The dashed lines represent the LSE-based linear regression model constructed for monometallic truncated octahedron NPs.
In this section, we discuss the universality of the LSE-based prediction model constructed in this study. As described in Section 3.1, due to the vastness of the chemical space for HEA NPs, we restricted both the model construction and the target NP shape to truncated octahedra. However, the LSE-based model can be universally applied to any HEA NP shape, provided the structure is given. This is because the universal NNP is capable of evaluating atomic energies for the given structure and thus computing the LSEs. The resulting LSEs can then be fed into the LSE-based model—originally constructed using monometallic data—to predict adsorption energies for the given HEA NP.
To verify the shape-independence of LSE-based prediction models, we evaluated its performance on two test sets: HEA146 in the form of a regular octahedron and HEA147 in the form of an icosahedron, as shown in the respective subsets in Figures 7(a) and (b). These shapes were selected because they are among the most prevalent geometries for fcc NPs. Figures 7(a) and (b) present the predicted CO adsorption energies from the NNP for all possible ontop sites on the regular octahedron and icosahedron, respectively. The corresponding LSE-based predictions are represented by the dashed lines as a function of LSE. For both NP shapes, we found clear correlations with the predictions obtained from the LSE-based model. The RMSE were 0.156 eV for the regular octahedron and 0.142 eV for the icosahedron, comparable to the RMSE of 0.150 eV achieved for truncated octahedra. These results demonstrate that our predictive framework exhibits universality with respect to NP shape within this scope.
It should be noted that while regular octahedra and icosahedra showcase the applicability of the LSE-based method for commonly explored fcc morphologies, many other potential NP shapes and surface reconstructions remain unexplored in this study. We focus here on the representative geometries to illustrate the applicability of the method, recognizing that a comprehensive proof of “universality” would require testing across more diverse crystal structures and potential defect sites of the HEA NPs. These aspects, along with extensions to other chemical compositions and surface features, are promising avenues for future research.
3.5 Comparison with GCN-based prediction model
Fig.  8: (a) Adsorption energies of CO calculated using the NNP for each on-top adsorption site of monometallic NPs $\mathrm{M}_{\mathrm{n}}$ with respect to the GCN values. Solid lines represent the linear regressions of the adsorption energies of a CO molecule at the on-top sites of each monometallic NP and the GCN values. (b) Adsorption energies of CO calculated using the NNP for each on-top adsorption site of 20 HEA201 NPs with respect to the GCN values. The dashed lines represent the same linear regressions as in (a). Circles, squares, and triangles at each datapoint represent the facet, edge, and corner CO adsorption sites, respectively.
In this section, we employ a GCN descriptor—a scalar descriptor similar to LSE—to construct a prediction model based on monometallic data and compare its predictive accuracy. The GCN descriptor quantifies the environment of the adsorption site by counting the coordination number of the nearest-neighbor atoms, and it is defined in Ref. 49 Following the same approach used for the LSE-based predictive model of Equation (4), we examined the correlation between CO adsorption energies on monometallic surfaces and GCN values. Linear regression models were then constructed for each element type. As shown in Figure 8(a), similar to LSE, a linear model accurately represents the relationship, achieving a precision of approximately 0.06 eV. The parameters of the linear regression models are summarized in Table 4 in Appendix section.
Subsequently, we applied these linear models to predict the adsorption energies of the 20 HEA NPs presented in the Results section to evaluate the prediction accuracy. As depicted in Figure 8(b), although the GCN-based predictions correlate with the directly computed adsorption energies, the discrete nature of GCN leads to variations of approximately from 1 to 1.5 eV even for identical GCN values. The resulting RMSE was 0.278 eV, approximately twice as large as that obtained using the LSE-based predictions, indicating that GCN alone is insufficient for capturing the environmental changes in multicomponent systems. This trend is consistent with findings reported by Namba et al. 51 Nonetheless, future improvements may be possible by combining GCN with other descriptors, such as LSE, to enhance the predictive capabilities of monometallic data-based adsorption energy prediction models.
@@ 4 Discussion
Finally, we discuss potential applications of the proposed method for designing chemical reactions for catalyst and device development. Previous studies examined the role of atomic energy in increasing the accuracy and efficiency of NNPs and validated atomic-energy mapping results 64, 68, 69, 70, 71. Recently, studies have attempted to gain chemical insight from atomic energies 72. However, to the best of our knowledge, the proposed LSE prediction method is the first to demonstrate that atomic energy can serve as a descriptor for the efficient evaluation of catalytic properties. Moreover, in contrast to atomic energies, which are absolute quantities, LSEs are relative quantities and are thus expected to be less sensitive to differences in computational methods, such as the treatment of basis functions and inner-shell electrons. The proposed approach enhances chemical reaction design—a crucial component of machine learning-based material design—which has garnered significant attention in the scientific community over the past few years due to its promise to accelerate the discovery and development of useful materials.73, 74.
@@ 5 Conclusion
Building upon this work, future research can extend the application of the LSE descriptor to other molecular species and reaction systems, validating and enhancing its predictive accuracy across a broader spectrum of catalytic processes. Integrating the LSE descriptor with advanced machine learning algorithms could facilitate large-scale screening of HEA compositions, accelerating the discovery of optimal catalysts for specific reactions. Moreover, combining the LSE descriptor with additional descriptors, such as surface microstructural features or GCN, offers the potential to further refine predictive accuracy by accounting for local atomic environments in more detail.49, 50, 51, 52, 11, 53, 54, 55, 56, 10
In this study, we assumed that the HEA NPs have random elemental configurations following a uniform distribution. However, short-range order is also important in high-entropy alloys75, 76, 77. By combining our method with techniques for quantifying short-range order using E(3)-equivariant graph neural networks 78, we can potentially obtain more accurate distributions of reactivity. This integration would allow us to capture the effects of atomic arrangements more precisely, leading to improved predictions of catalytic properties.
Moreover, this approach holds the potential to consider more realistic environmental conditions, including the behavior of catalysts in solution or under operational settings. Incorporating factors such as solvent effects, temperature, and pressure into the screening process would enhance the relevance and applicability of the predictions, leading to more effective and practical catalyst designs. Furthermore, applying the LSE framework to other material systems, such as high-entropy nitrides, oxides, and carbides, could further expand its impact on materials design.
@@ 6.1 Modeling
We modeled CO adsorption on the NPs, i.e., CO/$\mathrm{M}_{\mathrm{n}}$, using an atomic simulation environment (ASE) 79, 80. The initial lattice constants (LCs) of $\mathrm{M}_{\mathrm{n}}$ were determined via bulk calculations. For $\mathrm{HEA}_{\mathrm{201}}$, the largest bulk was used as the initial LC. A 15-Å vacuum region was inserted in all supercells to minimize cell-to-cell interactions. The initial structures of $\mathrm{CO/M}_{\mathrm{n}}$ and $\mathrm{CO/HEA}_{\mathrm{201}}$ were derived by placing CO on the on-top sites of the optimized structures of $\mathrm{M}_{\mathrm{n}}$ and $\mathrm{HEA}_{\mathrm{201}}$, respectively, with the distance between the C atom and the adsorption-site metal atom M set as 2 Å.
@@ 6.2 NNP calculations
The NNP used was M3GNet, a universal NNP proposed by Chen and Ong 36. This M3GNet NNP was trained on approximately 180,000 crystal environments at the PBE or PBE+U levels of theory from the Materials Project 81, covering 89 elements. The crystal structures of the bulk $\mathrm{M=Ir,Pd,Pt,Rh,Ru}$ were assumed to be fcc. Although $\mathrm{Ru}$ exhibits a hexagonal close-packed form at room temperature, $\mathrm{Ru}$ with an fcc structure can be created using NPs 82. The atomic energy of the bulk fcc metal M was determined using the energy corresponding to the minimum value obtained when varying the LC of the material at intervals of approximately 0.01 $\mathrm{\AA}$. The resulting atomic-energy values and corresponding LCs are presented in Table 2. The Broyden–Fletcher–Goldfarb–Shanno (BFGS) algorithm was used for structural optimizations, with a maximum step of 0.005 Å. In the CO/M and CO/HEA structural optimization calculations, the M and HEA structures were fixed as stable isolated systems. Only the CO and adsorption-site atoms were relaxed, with the constraint that the adsorbed molecules occupy the top site. Structural optimizations were performed until the force acting on each atom was 0.001 eV/Å.
@@ 6.3 DFT calculations
The DFT calculations were performed using the Vienna Ab initio Simulation Package (VASP), version 5.4.4 83, 84, 85. We employed the PBE generalized gradient approximation functional as the exchange-correlation functional86. The core electrons were treated using the projector-augmented wave method 87, 88. Electronic structures were optimized using the blocked Davidson iteration scheme within a spin-restricted approximation. The cutoff energy for the plane-wave functions was set as 400 eV. Atomic coordinates were optimized using a conjugate-gradient algorithm with a convergence threshold of 0.01 eV/Å.
6.4 GCN evaluations
The GCN descriptor is evaluated according to the following equation, which is identical to the one defined in Ref. 49
$\text{GCN}(i)=\sum_{j=1}^{n_{i}}\frac{cn(j)}{cn_{\max}}.$ (5)
Here, $i$ represents the index of the surface atom of the NPs. $n_{i}$ is the coordination number of the $i$-th atom, and $cn_{\max}$ is the coordination number of the $i$-th atom in its bulk environment, which is 12 for FCC metals considered in this study. $cn(j)$ represents the coordination number of each of the $n_{i}$ atoms coordinated to the $i$-th atom. In this study, GCN was evaluated for structures optimized using the universal NNP M3GNet, following the same procedure as for the evaluation of the LSE.
@@ 6.5 Nonlinear regressions
To capture the nonlinearity between LSE and CO adsorption energy, we trained a nonlinear regression model using all possible ontop adsorption sites (732 in total) from six randomly selected HEA nanoparticles out of the 20 used for validating the linear LSE-based model. For validation, all possible ontop adsorption sites on the remaining 14 HEA nanoparticles were utilized. KRRs were executed using scikit-learn version 1.2.2 89. A Gaussian kernel was selected as the kernel function of the KRR. The hyperparameters for each model were optimized over 100 iterations using a randomized search. QCL regressions were implemented using scikit-qulacs version 0.5.090. Qulacs version 0.5.6 was used as the quantum-circuit simulator91. The number of qubits in the QCL regression model was 4. The number of iterations of the parameterized variational quantum circuit of the model corresponding to the weights of the neural network was 6. The timestep for the time evolution operator in parameterized variational quantum circuits was set as 0.5. The BFGS algorithm was used to update the parameters of the QCL model.
@@
Table 3: Parameters $\alpha$ and $\beta$ of the regression lines between the CO adsorption energy and the local surface energy (LSE) for monometallic systems, obtained using NNP and DFT methods. The values correspond to Figures 3 and 10, respectively, where NNP and DFT in parentheses indicate the method used to obtain these parameters.
@@
Fig.  9: Correlation between the adsorption energies of a CO molecule on monometallic NPs calculated directly using the NNP and those predicted through regression. Circles, squares, and triangles at each datapoint represent the facet, edge, and corner CO adsorption sites, respectively.
Fig.  10: Correlation between adsorption energies of a CO molecule on monometallic NPs calculated directly via DFT and the LSEs corresponding to each adsorption site. The CO adsorption energies for the 38- and 79-atom NPs, as well as the datapoints corresponding to the hollow sites that converged on the Pd NPs, are shown as open markers, as they were not included in the linear regression data points. Circles, squares, and triangles at each datapoint represent the facet, edge, and corner CO adsorption sites, respectively.
Fig.  11: (a) Comparison and (b) correlation between the adsorption energy predicted via LSE and the adsorption energy of $\mathrm{CO/HEA}_{201}$ directly calculated using the NNP. We limit the comparison to 2440 sites, encompassing all on-top sites across the 20 structures. For each element (Ir, Pd, Pt, Rh, Ru) for 14 patterns of IrPdPtRhRu $\text{HEA}_{201}$ NPs, the energies are corrected using nonlinear regression models based on (c) KRR and (d) QCL regression. These values are plotted against values directly obtained using the NNP, which serve as the standard for accuracy.
@@
The data and code required to reproduce the figures and tables are publicly accessible on GitHub https://github.com/TShiotaSS/lse. The datasets for this study are available on Figshare at DOI: https://doi.org/10.6084/m9.figshare.26973409.v2. They contain structural and energetic information for both HEA and monometallic NPs. The datasets also include adsorption energy data from M3GNet and VASP calculations. The provided data includes relaxed structures and DFT calculations performed at the PBE level using VASP. Additionally, key descriptors such as GCN and LSE from M3GNet are included. We modified the code to extract the atomic energies from the pretrained M3GNet model on Github at https://github.com/materialsvirtuallab/m3gnet, which can be found at https://github.com/TShiotaSS/lse/tree/main/scripts/m3gnet_each_atom_energy. The implementation for QCL regression used in this study is available at https://github.com/Qulacs-Osaka/scikit-qulacs/tree/main.
@@
Here, we summarize the parameters of the regression model for adsorption-energy prediction using the LSE descriptor as well as the results of the model accuracy validation. Figure 9 shows the parity plot of the training data and prediction results for the LSE linear regression model presented in Figure 2. The RMSE for predicting the training data was 0.035 eV. Figure 10 presents the regression results for the linear regression model between the CO adsorption energies on monometallic NPs obtained via DFT calculations and the LSE. Two datapoints corresponding to the hollow sites that converged on the facets of Pd in 38- and 79-atom monometallic NPs were excluded from the regression model. For small NPs, the entire system tended to exhibit molecular characteristics upon adsorption, making it difficult for a simple regression model to capture these effects . However, for larger NPs, a linear relationship with the LSE, similar to that observed for the NNP, was confirmed. The parameters of the regression models constructed using the adsorption energies from both the NNP and DFT are presented in Table 3. Figure 11 presents a parity plot of the prediction results for the data not included in the training data for the nonlinear regression model of HEA NP adsorption energies evaluated directly using the NNP. The prediction accuracies were 0.0580 and 0.0579 eV, respectively, indicating that, as demonstrated in this study, the regression models using KRR and QCL had comparable accuracy. Finally, the regression parameters of the GCN-based prediction model are summarized in Table 4.
Table 4: Parameters $\alpha$ and $\beta$ of the regression lines between the CO adsorption energy and the generalized coordination number (GCN) for monometallic systems. The values correspond to the lines in Figures 8(a) and (b)
Ir Pd Pt Rh Ru
$\alpha$(NNP) 0.185 0.139 0.173 0.078 0.121
$\beta$(NNP) -2.766 -2.403 -1.073 -1.946 -2.371